# Editable Concept Bottleneck Models

### v1

@@ 1 Introduction
Modern deep learning models, such as large language models zhao2023survey ; yang2024moral ; yang2024human ; xu2023llm ; yang2024dialectical and large multimodal yin2023survey ; ali2024prompt ; cheng2024multi , often exhibit intricate non-linear architectures, posing challenges for end-users seeking to comprehend and trust their decisions. This lack of interpretability presents a significant barrier to adoption, particularly in critical domains such as healthcare ahmad2018interpretable ; yu2018artificial and finance cao2022ai , where transparency is paramount. To address this demand, explainable artificial intelligence (XAI) models das2020opportunities ; hu2023seat ; hu2023improving have emerged, offering explanations for their behavior and insights into their internal mechanisms. Among these, Concept Bottleneck Models (CBMs) koh2020concept have gained prominence for explaining the prediction process of end-to-end AI models. In CBMs, they add a bottleneck layer for placing human-understandable concepts. In the prediction process, CBMs first predict the concept labels using the original input and then predict the final classification label using the predicted concept in the bottleneck layer, which provides a self-explained decision to users.
@@
Existing research on CBMs predominantly addresses two primary concerns: Firstly, CBMs heavily rely on laborious dataset annotation. Researchers have explored solutions to these challenges in unlabeled settings oikarinen2023label ; yuksekgonul2022post ; lai2023faithful . Secondly, the performance of CBMs often lags behind that of original models lacking the concept bottleneck layer, attributed to incomplete information extraction from original data to bottleneck features. Researchers aim to bridge this utility gap chauhan2023interactive . However, few of them considered the adaptivity or editability of CBMs, crucial aspects encompassing annotation errors, data privacy considerations, or concept updates. Actually, these demands are increasingly pertinent in the era of large models. We delineate the editable setting into three key aspects (illustrated in Figure 1):
@@
  • Concept-level: In CBMs, the concept set is pre-defined by LLMs or experts. However, in many cases, evolving situations demand concept updates, as evidenced by discoveries such as chronic obstructive pulmonary disease as a risk factor for lung cancer, and doctors have the requirements to add related concepts. For another example, recent research found a new factor, obesity sattar2020obesity ) are risky for severe COVID-19 and factors (e.g., older age, male gender, Asian race) are risk associated with COVID-19 infection rozenfeld2020model . On the other hand, one may also want to remove some spurious or unrelated concepts for the task. This demand is even more urgent in some rapidly evolving domains like the pandemic.
@@
We propose Editable Concept Bottleneck Models (ECBMs) to tackle these challenges. Specifically, compared to retraining, ECBMs provide a mathematically rigorous closed-form approximation for the above three settings to address editability within CBMs efficiently. Leveraging the influence function cook2000detection ; cook1980characterizations , we quantify the impact of individual data points, individual concept labels, and the concept for all data on model parameters. Despite the growing attention and utility of influence functions in machine learning koh2017understanding , their application in CBMs remains largely unexplored due to their composite structure, i.e., the intermediate representation layer. To the best of our knowledge, we are the first to work to fill this gap by demonstrating the effectiveness of influence functions in elucidating the behavior of CBMs, especially in identifying mislabeled data and discerning the data influence. Comprehensive experiments on benchmark datasets show that our ECBMs are efficient and effective. Our contributions are summarized as follows.
@@
  • To make CBMs able to remove data or concept influence without retraining, we propose the Editable Concept Bottleneck Models (ECBMs). Our approach in ECBMs offers a mathematically rigorous closed-form approximation. Furthermore, to improve computational efficiency, we present streamlined versions integrating EK-FAC.
@@ 2 Related Work
Concept Bottleneck Models. CBM koh2020concept stands out as an innovative deep-learning approach for image classification and visual reasoning. It introduces a concept bottleneck layer into deep neural networks, enhancing model generalization and interpretability by learning specific concepts. However, CBM faces two primary challenges: its performance often lags behind that of original models lacking the concept bottleneck layer, attributed to incomplete information extraction from the original data to bottleneck features. Additionally, CBM relies on laborious dataset annotation. Researchers have explored solutions to these challenges. chauhan2023interactive extend CBM into interactive prediction settings, introducing an interaction policy to determine which concepts to label, thereby improving final predictions. oikarinen2023label address CBM limitations and propose a novel framework called Label-free CBM. This innovative approach enables the transformation of any neural network into an interpretable CBM without requiring labeled concept data, all while maintaining high accuracy. Post-hoc Concept Bottleneck models yuksekgonul2022post can be applied to various neural networks without compromising model performance, preserving interpretability advantages. CBMs work on the image field also includes havasi2022addressing ; kim2023probabilistic ; keser2023interpretable ; sawada2022concept ; sheth2023auxiliary . Despite many works on CBMs, we are the first to investigate the interactive influence between concepts through influence functions. Our research endeavors to bridge this gap by utilizing influence functions in CBMs, thereby deciphering the interaction of concept models and providing an adaptive solution to concept editing. More related work refer to Appendix G.
@@ 3 Preliminaries
Concept Bottleneck Models. In this paper, we consider the original CBM, and we adopt the notations used by koh2020concept . We consider a classification task with a concept set denoted as $\bm{c}=\{p_{1},\cdots,p_{k}\}$ with each $p_{i}$ is a concept given by experts or LLMs, and a training dataset represented as $\mathcal{D}=\{z_{i}\}_{i=1}^{n}$, where $z_{i}=(x_{i},y_{i},c_{i})$. Here, for $i\in[n]$, $x_{i}\in\mathbb{R}^{m}$ represents the feature vector, $y_{i}\in\mathbb{R}^{d_{z}}$ denotes the label (with $d_{z}$ corresponding to the number of classes), and $c_{i}=(c_{i}^{1},\cdots,c_{i}^{k})\in\mathbb{R}^{k}$ represents the concept vector. In this context, $c_{i}^{j}$ represents the weight of the concept $p_{j}$. In CBMs, our goal is to learn two representations: one called concept predictor that transforms the input space to the concept space, denoted as $g:\mathbb{R}^{m}\to\mathbb{R}^{k}$, and another called label predictor that maps the concept space to the prediction space, denoted as $f:\mathbb{R}^{k}\to\mathbb{R}^{d_{z}}$. Usually, here the map $f$ is linear. For each training sample $z_{i}=(x_{i},y_{i},c_{i})$, we consider two empirical loss functions: one is from the input space to concept space, and the other is from concept space to output space:
@@
Influence function. The influence function is a measure of the dependence of the estimator on the value of any one of the points in the sample. Consider a neural network $\hat{\theta}=\arg\min_{\theta}\sum_{i=1}^{n}\ell(z_{i},\theta)$ with loss function $\ell$ and dataset $D=\{z_{i}\}_{i=1}^{n}$. Removing a point $z_{m}$, the model change is denoted as $\hat{\theta}_{-z_{m}}-\hat{\theta}$, where $\hat{\theta}_{-z_{m}}\triangleq\operatorname*{arg\,min}_{\theta}\sum_{i\neq m}\ell\left(z_{i},\theta\right)$. Influence function gives us an efficient approximation for $\hat{\theta}_{-z_{m}}-\hat{\theta}$ by parameterizing the training set using the weight $\epsilon$ assigned to the new sample and observing consequent changes in the optimal parameters $\hat{\theta}_{-z_{m}}(\epsilon)=\operatorname*{arg\,min}_{\theta}\sum_{i\neq m}\ell(z_{i},\theta)+\epsilon\cdot\ell(z_{m},\theta)$. Subsequently, our focus shifts to elucidating the influence of the sample $z_{m}$ on the parameter ${\theta}$. The response function is defined as the first-order Taylor approximation evaluated at $\epsilon=0$ to capture this effect, that is $\mathcal{I}_{\hat{\theta}}\left(z_{m}\right)\triangleq\left.\frac{\mathrm{d}{\hat{\theta}}}{\mathrm{d}\epsilon}\right|_{\epsilon=0}=-H^{-1}_{\hat{\theta}}\cdot\nabla_{\theta}\ell(z_{m},{\hat{\theta}})$, where $H^{-1}_{{\hat{\theta}}}=\nabla^{2}_{\theta}\sum_{i=1}^{n}\ell(z_{i},{\hat{\theta}})$ is the Hessian matrix of the loss function on the original training dataset with respect to ${\hat{\theta}}$. Then, a quadratic approximation is created for the empirical risk centered at $\hat{\theta}$. By taking a single Newton step with a step length of 1, we can estimate the change in parameters resulting from the exclusion of $z_{m}$ without the need to retrain the model. This estimation is obtained by computing $\hat{\theta}_{-z_{m}}-\hat{\theta}\approx-\mathcal{I}_{{\hat{\theta}}}(z_{m})$.
When the loss function $\ell$ is twice-differentiable and strongly convex in $\theta$, the Hessian $H_{{\hat{\theta}}}$ is positive definite and thus the influence function is well-defined. For non-convex loss, bartlett1953approximate proposed that the Hessian $H_{{\hat{\theta}}}$ can be replaced by $\hat{H}=G_{\hat{\theta}}+\lambda I$ where $G_{\hat{\theta}}$ is the Fisher information matrix defined by $n^{-1}\sum_{i=1}^{n}\nabla_{\theta}\ell(z_{i}.\theta)\nabla_{\theta}\ell(z_{i}.\theta)^{\mathrm{T}}$, $\lambda$ is a small damping term used to ensure the positive definiteness of $\hat{H}$. We can employ the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) method to further accelerate the computation. See Appendix A for details.
@@ 4.1 Concept Label-level Editable CBM
For simple neural networks, we can use the influence function approach directly to estimate the retrained model. However, for CBM architecture, if we intervene with the true concepts, the concept predictor $\hat{g}$ fluctuates to $\hat{g}_{e}$ accordingly. Observing that the input data of the label predictor is the output of the concept predictor, which is also changing. Therefore, we need to adopt a two-stage edition approach. Here we consider the influence function for (3) and (4) separately. We first edit the concept predictor from $\hat{g}$ to $\bar{g}_{e}$, and then edit from $\hat{f}$ to $\bar{f}_{e}$ based on our approximated concept predictor, by the following two theorems.
@@
The retrained label predictor $\hat{f}_{e}$ defined by (4) can be approximated by:
@@ 4.2 Concept-level Editable CBM
In this case, a set of concepts is removed due to incorrect attribution or spurious concepts, termed concept-level edit.  [^1] For convenience, in this paper, we only consider concept removal; our method can directly extend to concept insertion.Specifically, for the concept set $\bm{c}$, denote the erroneous concepts as $\{p_{r}\}$, $r\in M\subset[k]$, we aim to delete their concept labels in all training samples $\mathcal{D}$. We aim to investigate the impact of updating the concept set within the training data on the model’s predictions. It is notable that compared to the above concept label case, the dimension of output (input) of the retrained concept predictor (label predictor) will change. If we delete $t$ concepts from the dataset, then ${g}$ becomes ${g}^{\prime}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{k-t}$ and ${f}$ becomes ${f}^{\prime}:\mathbb{R}^{k-t}\rightarrow\mathbb{R}^{d_{z}}$. More specifically, if we retrain the CBM with the revised dataset, the corresponding concept predictor becomes:
@@
To overcome the challenge, our strategy is to develop some transformations that need to be performed on $\hat{g}_{-p_{M}}$ to align its dimension with $\hat{g}$ so that we can apply the influence function to edit the CBM. We achieve this by mapping $\hat{g}_{-p_{M}}$ to $\mathrm{P}(\hat{g})=\hat{g}^{*}_{-p_{M}}$ via some mapping $\mathrm{P}$, which has the same amount of parameters as $\hat{g}$ and has the same predicted concepts $\hat{g}^{*j}_{-p_{M}}$, as $\hat{g}^{j}_{-p_{M}}$ for all $j\in[k]-M$. We achieve this effect by inserting a zero row vector into the $r$-th row of the matrix in the final layer of $\hat{g}_{-p_{M}}$ for $r\in M$. Thus, we can see that the mapping $P$ is one-to-one.
Moreover, assume the parameter space of $\hat{g}$ is $\Gamma$ and denote the $t$-th layer parameter matrix of $\hat{g}$ as $W^{(t)}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, $t\in[L]$. We have $\hat{g}^{*}_{-p_{M}}\in\Gamma_{0}=\{W^{(L)}_{r,:}=0,r\in M\}\subseteq\Gamma$ and $\hat{g}^{*}_{-p_{M}}$ is the optimal model of the following objective function:
$\hat{g}^{*}_{-p_{M}}=\operatorname*{arg\,min}_{g^{\prime}\in\Gamma_{0}}\sum_{j\notin M}^{k}\sum_{i=1}^{n}L_{C_{j}}(g^{\prime j}(x_{i}),c_{i}^{j}),$ (7)
i.e., it is the optimal model of the concept predictor loss on the remaining concepts under the constraint $\Gamma_{0}$. Now we can apply the influence function to edit $\hat{g}$ to approximate $\hat{g}^{*}_{-p_{M}}$ with the restriction on the value of 00 for rows indexed by $M$ with the last layer of the neural network, denoted as $\bar{g}^{*}_{-p_{M}}$. After that, we remove from $\bar{g}^{*}_{-p_{M}}$ the parameters initially inserted to fill in the dimensional difference, which always equals 00 because of the restriction we applied in the editing stage, thus approximating the true edited concept predictor $\hat{g}_{-p_{M}}$. We now detail the editing process from $\hat{g}$ to $\hat{g}^{*}_{-p_{M}}$ using the following theorem.
@@
For the retrained concept predictor $\hat{g}_{-p_{M}}$ defined in (6), we map it to $\hat{g}^{*}_{-p_{M}}$ as (7). And we can edit the initial $\hat{g}$ to $\hat{g}^{*}_{-p_{M}}$ as:
@@
For the revised retrained label predictor $\hat{f}_{p_{M}=0}$ defined by (8), we can edit the initial label predictor $\hat{f}$ to $\bar{f}_{p_{M}=0}$ by the following equation as a substitute for $\hat{f}_{p_{M}=0}$:

### v2

@@ 1 Introduction
Modern deep learning models, such as large language models (Zhao et al., 2023; Yang et al., 2024a; b; Xu et al., 2023; Yang et al., 2024c) and large multimodal (Yin et al., 2023; Ali et al., 2024; Cheng et al., 2024), often exhibit intricate non-linear architectures, posing challenges for end-users seeking to comprehend and trust their decisions. This lack of interpretability presents a significant barrier to adoption, particularly in critical domains such as healthcare (Ahmad et al., 2018; Yu et al., 2018) and finance (Cao, 2022), where transparency is paramount. To address this demand, explainable artificial intelligence (XAI) models (Das & Rad, 2020; Hu et al., 2023b; a) have emerged, offering explanations for their behavior and insights into their internal mechanisms. Among these, Concept Bottleneck Models (CBMs) (Koh et al., 2020) have gained prominence for explaining the prediction process of end-to-end AI models. CBMs add a bottleneck layer for placing human-understandable concepts. In the prediction process, CBMs first predict the concept labels using the original input and then predict the final classification label using the predicted concept in the bottleneck layer, which provides a self-explained decision to users.
@@
Existing research on CBMs predominantly addresses two primary concerns: Firstly, CBMs heavily rely on laborious dataset annotation. Researchers have explored solutions to these challenges in unlabeled settings (Oikarinen et al., 2023; Yuksekgonul et al., 2023; Lai et al., 2023). Secondly, the performance of CBMs often lags behind that of original models lacking the concept bottleneck layer, attributed to incomplete information extraction from original data to bottleneck features. Researchers aim to bridge this utility gap (Sheth & Ebrahimi Kahou, 2023; Yuksekgonul et al., 2023; Espinosa Zarlenga et al., 2022). However, few of them considered the adaptivity or editability of CBMs, crucial aspects encompassing annotation errors, data privacy considerations, or concept updates. Actually, these demands are increasingly pertinent in the era of large models. We delineate the editable setting into three key aspects (illustrated in Figure 1):
@@
  • Concept-level: In CBMs, the concept set is pre-defined by LLMs or experts. However, in many cases, evolving situations demand concept updates, as evidenced by discoveries such as chronic obstructive pulmonary disease as a risk factor for lung cancer, and doctors have the requirements to add related concepts. For another example, recent research found a new factor, obesity (Sattar et al., 2020) are risky for severe COVID-19 and factors (e.g., older age, male gender, Asian race) are risk associated with COVID-19 infection (Rozenfeld et al., 2020). On the other hand, one may also want to remove some spurious or unrelated concepts for the task. This demand is even more urgent in some rapidly evolving domains like the pandemic.
@@
We propose Editable Concept Bottleneck Models (ECBMs) to tackle these challenges. Specifically, compared to retraining, ECBMs provide a mathematically rigorous closed-form approximation for the above three settings to address editability within CBMs efficiently. Leveraging the influence function (Cook, 2000; Cook & Weisberg, 1980), we quantify the impact of individual data points, individual concept labels, and the concept for all data on model parameters. Despite the growing attention and utility of influence functions in machine learning (Koh & Liang, 2017), their application in CBMs remains largely unexplored due to their composite structure, i.e., the intermediate representation layer.
To the best of our knowledge, we are the first to work to fill this gap by demonstrating the effectiveness of influence functions in elucidating the behavior of CBMs, especially in identifying mislabeled data and discerning the data influence. Comprehensive experiments on benchmark datasets show that our ECBMs are efficient and effective. Our contributions are summarized as follows.
@@
  • To make CBMs able to remove data or concept influence without retraining, we propose the Editable Concept Bottleneck Models (ECBMs). Our approach in ECBMs offers a mathematically rigorous closed-form approximation. Furthermore, to improve computational efficiency, we present streamlined versions integrating Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC).
@@ 2 Related Work
Concept Bottleneck Models. CBM (Koh et al., 2020) stands out as an innovative deep-learning approach for image classification and visual reasoning. It introduces a concept bottleneck layer into deep neural networks, enhancing model generalization and interpretability by learning specific concepts. However, CBM faces two primary challenges: its performance often lags behind that of original models lacking the concept bottleneck layer, attributed to incomplete information extraction from the original data to bottleneck features. Additionally, CBM relies on laborious dataset annotation. Researchers have explored solutions to these challenges. Chauhan et al. (2023) extend CBM into interactive prediction settings, introducing an interaction policy to determine which concepts to label, thereby improving final predictions. Oikarinen et al. (2023) address CBM limitations and propose a novel framework called Label-free CBM. This innovative approach enables the transformation of any neural network into an interpretable CBM without requiring labeled concept data, all while maintaining high accuracy. Post-hoc Concept Bottleneck models (Yuksekgonul et al., 2023) can be applied to various neural networks without compromising model performance, preserving interpretability advantages. CBMs work on the image field also includes (Havasi et al., 2022; Kim et al., 2023; Keser et al., 2023; Sawada & Nakamura, 2022; Sheth & Kahou, 2023). Despite many works on CBMs, we are the first to investigate the interactive influence between concepts through influence functions. Our research endeavors to bridge this gap by utilizing influence functions in CBMs, thereby deciphering the interaction of concept models and providing an adaptive solution to concept editing. For more related work, please refer to Appendix H.
@@ 3 Preliminaries
Concept Bottleneck Models. In this paper, we consider the original CBM, and we adopt the notations used by Koh et al. (2020). We consider a classification task with a concept set denoted as $\bm{c}=\{p_{1},\cdots,p_{k}\}$ with each $p_{i}$ is a concept given by experts or LLMs, and a training dataset represented as $\mathcal{D}=\{z_{i}\}_{i=1}^{n}$, where $z_{i}=(x_{i},y_{i},c_{i})$. Here, for $i\in[n]$, $x_{i}\in\mathbb{R}^{m}$ represents the feature vector, $y_{i}\in\mathbb{R}^{d_{z}}$ denotes the label (with $d_{z}$ corresponding to the number of classes), and $c_{i}=(c_{i}^{1},\cdots,c_{i}^{k})\in\mathbb{R}^{k}$ represents the concept vector. In this context, $c_{i}^{j}$ represents the weight of the concept $p_{j}$. In CBMs, our goal is to learn two representations: one called concept predictor that transforms the input space to the concept space, denoted as $g:\mathbb{R}^{m}\to\mathbb{R}^{k}$, and another called label predictor that maps the concept space to the prediction space, denoted as $f:\mathbb{R}^{k}\to\mathbb{R}^{d_{z}}$. Usually, here the map $f$ is linear. For each training sample $z_{i}=(x_{i},y_{i},c_{i})$, we consider two empirical loss functions: one is from the input space to concept space, and the other is from concept space to output space:
@@
Influence function. The influence function is a measure of the dependence of the estimator on the value of any one of the points in the sample. Consider a neural network $\hat{\theta}=\arg\min_{\theta}\sum_{i=1}^{n}\ell(z_{i};\theta)$ with loss function $\ell$ and dataset $D=\{z_{i}\}_{i=1}^{n}$. If we remove a point $z_{m}$ from the training dataset, the parameters become $\hat{\theta}_{-z_{m}}=\arg\min_{\theta}\sum_{i\neq m}\ell(z_{i};\theta)$. Influence function gives us an efficient approximation for $\hat{\theta}_{-z_{m}}$ by defining a response function as $\hat{\theta}_{\epsilon,-z_{m}}=\underset{\theta\in\Theta}{\operatorname{argmin}}\frac{1}{n}\sum_{i=1}^{n}L(z_{i};\theta)+\epsilon L(z_{m};\theta).$ Perform first-order Taylor expansion on the gradient of the objective function corresponding to the $\operatorname*{arg\,min}$ process of the response function, we can obtain the influence function defined by
$\mathcal{I}_{\hat{\theta}}\left(z_{m}\right)\triangleq\left.\frac{\mathrm{d}{\hat{\theta}}}{\mathrm{d}\epsilon}\right|_{\epsilon=0}=-H^{-1}_{\hat{\theta}}\cdot\nabla_{\theta}\ell(z_{m};{\hat{\theta}}),$
which can evaluate the influence of $z_{m}$ on the parameters. When the loss function $\ell$ is twice-differentiable and strongly convex in $\theta$, the Hessian $H_{{\hat{\theta}}}$ is positive definite and thus the influence function is well-defined. For non-convex loss, Bartlett (1953) proposed that the Hessian $H_{{\hat{\theta}}}$ can be replaced by $\hat{H}=G_{\hat{\theta}}+\lambda I$ where $G_{\hat{\theta}}$ is the Fisher information matrix defined by $n^{-1}\sum_{i=1}^{n}\nabla_{\theta}\ell(z_{i};\theta)\nabla_{\theta}\ell(z_{i};\theta)^{\mathrm{T}}$, $\lambda$ is a small damping term used to ensure the positive definiteness of $\hat{H}$. We can employ the Eigenvalue-corrected Kronecker-Factored Approximate Curvature (EK-FAC) method to further accelerate the computation. See Appendix B for details.
@@ 4.1 Concept Label-level Editable CBM
For simple neural networks, we can use the influence function approach directly to estimate the retrained model. However, for CBM architecture, if we intervene with the true concepts, the concept predictor $\hat{g}$ fluctuates to $\hat{g}_{e}$ accordingly. Observing that the input data of the label predictor is the output of the concept predictor, which is also changing. Therefore, we need to adopt a two-stage edition approach. Here we consider the influence function for equation 3 and equation 4 separately. We first edit the concept predictor from $\hat{g}$ to $\bar{g}_{e}$, and then edit from $\hat{f}$ to $\bar{f}_{e}$ based on our approximated concept predictor, by the following two theorems.
@@
The retrained label predictor $\hat{f}_{e}$ defined by equation 4 can be approximated by:
@@ 4.2 Concept-level Editable CBM
In this case, a set of concepts is removed due to incorrect attribution or spurious concepts, termed concept-level edit.  [^1] For convenience, in this paper, we only consider concept removal; our method can directly extend to concept insertion.Specifically, for the concept set, denote the erroneous concept index set as $M\subset[k]$, we aim to delete these concept labels in all training samples. We aim to investigate the impact of updating the concept set within the training data on the model’s predictions. It is notable that compared to the above concept label case, the dimension of output (input) of the retrained concept predictor (label predictor) will change. If we delete $t$ concepts from the dataset, then ${g}$ becomes ${g}^{\prime}:\mathbb{R}^{m}\rightarrow\mathbb{R}^{k-t}$ and ${f}$ becomes ${f}^{\prime}:\mathbb{R}^{k-t}\rightarrow\mathbb{R}^{d_{z}}$. More specifically, if we retrain the CBM with the revised dataset, the corresponding concept predictor becomes:
@@
To overcome the challenge, our strategy is to develop some transformations that need to be performed on $\hat{g}_{-p_{M}}$ to align its dimension with $\hat{g}$ so that we can apply the influence function to edit the CBM. We achieve this by mapping $\hat{g}_{-p_{M}}$ to $\hat{g}^{*}_{-p_{M}}\triangleq\mathrm{P}(\hat{g}_{-p_{M}})$, which has the same amount of parameters as $\hat{g}$ and has the same predicted concepts $\hat{g}^{*}_{-p_{M}}(j)$ as $\hat{g}_{-p_{M}}(j)$ for all $j\in[k]-M$. We achieve this effect by inserting a zero row vector into the $r$-th row of the matrix in the final layer of $\hat{g}_{-p_{M}}$ for $r\in M$. Thus, we can see that the mapping $P$ is one-to-one. Moreover, assume the parameter space of $\hat{g}$ is $T$ and that of $\hat{g}^{*}_{-p_{M}}$, $T_{0}$ is the subset of is the subset of $T$. Noting that $\hat{g}^{*}_{-p_{M}}$ is the optimal model of the following objective function:
$\hat{g}^{*}_{-p_{M}}=\operatorname*{arg\,min}_{g^{\prime}\in T_{0}}\sum_{j\notin M}^{k}\sum_{i=1}^{n}L_{C_{j}}(g^{\prime j}(x_{i}),c_{i}^{j}),$ (7)
i.e., it is the optimal model of the concept predictor loss on the remaining concepts under the constraint $T_{0}$. Now we can apply the influence function to edit $\hat{g}$ to approximate $\hat{g}^{*}_{-p_{M}}$ with the restriction on the value of 00 for rows indexed by MM with the last layer of the neural network, denoted as $\bar{g}^{*}_{-p_{M}}$. After that, we remove from $\bar{g}^{*}_{-p_{M}}$ the parameters initially inserted to fill in the dimensional difference, which always equals 00 because of the restriction we applied in the editing stage, thus approximating the true edited concept predictor $\hat{g}_{-p_{M}}$. We now detail the editing process from $\hat{g}$ to $\hat{g}^{*}_{-p_{M}}$ using the following theorem.
@@
For the retrained concept predictor $\hat{g}_{-p_{M}}$ defined in equation 6, we map it to $\hat{g}^{*}_{-p_{M}}$ as equation 7. And we can edit the initial $\hat{g}$ to $\hat{g}^{*}_{-p_{M}}$ as:
@@
For the revised retrained label predictor $\hat{f}_{p_{M}=0}$ defined by equation 8, we can edit the initial label predictor $\hat{f}$ to $\bar{f}_{p_{M}=0}$ by the following equation as a substitute for $\hat{f}_{p_{M}=0}$: